# A stringy dispersion relation for field theory

### v2

@@ 1 Introduction
where $(x)_{n}$ is the Pochammer symbol. The representation is independent of the parameter $\lambda$ and converges when $\text{Re}(\lambda)>0$. We see that it matches the string field theory expectations outlined above. It [^4] Disclaimer: It is not obvious how the parameter is exactly related to that in string field theory.. This motivates the search for such parametric representations for tree-level string amplitudes in general.
@@ 2 The two-channel symmetric dispersion relation
is the $s$-channel discontinuity of the amplitude across the contour $\mathcal{C}_{1}$. The dispersion relation is independent of the parameter $\lambda$. Further, when we plug in $\lambda=-t$, we recover the fixed-$t$ dispersion relation, while $\lambda=-s$ gives the fixed-$s$. Thus, we see that the parameter precisely plays the role of the “elasticity” parameter of the string worldsheet mentioned before (see Fig. (1)), which enables us to deform between and equate the $s$-channel and $t$-channel representations. This exhibits the “worldsheet duality” in string theory very naturally, prompting the name ”stringy” for such a dispersion relation. Let us now see how to derive it.
@@ 2.1 Derivation
We assume that the amplitude satisfies $\lim_{|s|\to\infty}\left|\mathcal{M}\left(s,t\right)\right|=0$ in some domain $\mathcal{D}$ of $t$. Let’s consider an arbitrary, as-of-now undetermined function $\hat{s}_{2}^{(\lambda)}(\sigma,s,t)$. For now, we only demand that $\hat{s}_{2}^{(\lambda)}(\sigma,s,t)$ be $s\leftrightarrow t$ symmetric and $\lim_{|\sigma|\to\infty}\hat{s}_{2}^{(\lambda)}(\sigma,s,t)\in\mathcal{D}$. Then we have
@@
This ensures that the following contour integral along $\Gamma$, a large circle at infinity, vanishes.
If the amplitude grows faster at large $\sigma$ than we have assumed, we can introduce additional counter-terms/subtraction-terms in the integrand to ensure that the integral vanishes. We discuss such cases in the next section.
Now, let us contract the contour. We pick up singularities from both the kernel and the amplitude. The contributions from the amplitude lead to the following.
@@
where $\mathcal{A}^{(s)}(s,t)$ is the discontinuity of the amplitude in the $s$-channel along the contour $\mathcal{C}_{1}$, and $\mathcal{A}^{(t)}(s,t)$ is the discontinuity of the amplitude in the $t$-channel along the contour $\mathcal{C}_{2}$. Due to $s\leftrightarrow t$ symmetry of the amplitude, $\mathcal{A}^{(t)}(s,t)=\mathcal{A}^{(s)}(t,s)$. We now demand that the two terms in (2.6) be equal. Let us do a coordinate transformation $\sigma=\sigma(\tau)$ in the second term. Then we get the following.
$\frac{1}{\pi}\int\mathrm{d}\tau\frac{d\sigma}{d\tau}\mathcal{H}^{(\lambda)}(\sigma(\tau),s,t)\mathcal{A}^{(t)}(\sigma(\tau),\hat{s}_{2}^{(\lambda)}(\sigma(\tau),s,t))=\frac{1}{\pi}\int\mathrm{d}\tau\frac{d\sigma}{d\tau}\mathcal{H}^{(\lambda)}(\sigma(\tau),s,t)\mathcal{A}^{(s)}(\hat{s}_{2}^{(\lambda)}(\sigma(\tau),s,t),\sigma(\tau))\,.$ (2.7)
Equality of the two terms in (2.6) requires
@@
$\frac{d}{d\sigma}\ln\frac{(\sigma-s)(\sigma-t)}{(\lambda+\sigma)}=\frac{d}{d\tau}\ln\frac{(\tau-s)(\tau-t)}{(\lambda+\tau)}\,$, (2.9)
@@
where $c$ is an integration constant. The second condition fixes $c=1$. We thus get a quadratic equation for $\tau$ with two solutions:
@@
The first solution has no $s,t$ dependence and therefore we discard it. Using the second solution, the contribution from the singularities of the amplitude is
@@
Now, we come to the singularities that arise from the kernel $\mathcal{H}(s,t,\lambda)$. Let us first discuss the singularity at $\sigma=-\lambda$. Due to our assumed large $\sigma$ behaviour, we have $\left|\mathcal{M}\left(\sigma,-\lambda\right)\right|\underset{\sigma\to\infty}{\sim}0$, which by two-channel symmetry leads to $\left|\mathcal{M}\left(-\lambda,\sigma\right)\right|\underset{\sigma\to\infty}{\sim}0$ and therefore,
@@
Figure 2: A typical depiction of the singularities in the integrand in the complex-$\sigma$ plane. Here $s=1+i,\>t=2+i,\>\lambda=1$ and there is an s-channel cut for $s>2$ and its t-channel counterpart. In the integrand, the t-channel cut counterpart gets mapped to the line joining $(1/3,7/3)$ and (1,0)(-1,0).
@@ 3.1 Derivation
when $\lim_{|\sigma|\to\infty}\hat{s}_{2}^{(\lambda)}(\sigma,s_{1},s_{2},s_{3})\in\mathcal{D}$. $\hat{s}_{2}^{(\lambda)}(\sigma,s_{1},s_{2},s_{3})$ is as-of-yet an undetermined function which we demand to be fully crossing symmetric. The following contour integral along a large circle at infinity vanishes.
@@ 4.3 One-loop box diagram
The one-loop box diagram in four dimensions with two massive and two massless propagators takes the following form [33]
@@ 5.4 Bounds on weakly-coupled gravitational EFTs
Figure 8: Bounds with gravity for dilaton-like amplitudes in $d=5$. The region to the right of the curves is allowed. As can be seen, the bounds converge as $k_{max}$ is increased. We define $W_{1,0}=-\dfrac{2g_{2}}{8\pi G}$ and $W_{0,1}=\dfrac{g_{3}}{8\pi G}$ to match the conventions in [49] and aid comparison.

### v3

@@ 1 Introduction
where $(x)_{n}$ is the Pochammer symbol. The representation is independent of the parameter $\lambda$ and converges when $\text{Re}(\lambda)>0$. We see that it matches the string field theory expectations outlined above [^4] Disclaimer: It is not obvious how the parameter is exactly related to that in string field theory.. This motivates the search for such parametric representations for tree-level string amplitudes in general.
@@ 2 The two-channel symmetric dispersion relation
is the $s$-channel discontinuity of the amplitude across the contour $\mathcal{C}_{1}$. The dispersion relation is independent of the parameter $\lambda$. Further, when we plug in $\lambda=-t$, we recover the fixed-$t$ dispersion relation, while $\lambda=-s$ gives the fixed-$s$. Thus, we see that the parameter precisely plays the role of the “elasticity” parameter of the string worldsheet mentioned before (see Fig. (1)), which enables us to deform between and equate the $s$-channel and $t$-channel representations. This exhibits the “worldsheet duality” in string theory very naturally, prompting the name “stringy” for such a dispersion relation. Let us now see how to derive it.
@@ 2.1 Derivation
Assume that the amplitude satisfies $\lim_{|s|\to\infty}\left|\mathcal{M}\left(s,t\right)\right|=0$ in some domain $\mathcal{D}$ of $t$. Then the following contour integral along $\Gamma$, a large circle at infinity, vanishes for $t\in\mathcal{D}$.
$\frac{1}{2\pi i}\int_{\Gamma}\mathrm{d}\sigma\frac{1}{\sigma-s}\mathcal{M}\left(\sigma,t\right)=0$ (2.4)
Shrinking the contour, we get a contribution from the pole at $\sigma=s$ which is $\mathcal{M}(s,t)$ and another contribution from the singularities of the amplitude, given in terms of the $s$-channel discontinuity $\mathcal{A}^{(s)}(s,t)$ along the contour $\mathcal{C}_{1}$. This leads to the fixed-$t$ dispersion relation.
$\mathcal{M}(s,t)=\frac{1}{\pi}\int_{\mathcal{C}_{1}}d\sigma\frac{1}{\sigma-s}\mathcal{A}(\sigma,t)$ (2.5)
However, since the result in (2.4) only depends on the large $s$ behaviour of the amplitude, we can actually do a more general analysis. We can replace $\mathcal{M}\left(\sigma,t\right)$ by $\mathcal{M}\left(\hat{s}_{1}(\sigma,s,t),\hat{s}_{2}^{(\lambda)}(\sigma,s,t)\right)$ and as long as when $\sigma\to\infty$, $\hat{s}_{1}(\sigma,s,t)\to\infty$ and $\hat{s}_{2}^{(\lambda)}(\sigma,s,t)\to-\lambda$ for some $\lambda\in\mathcal{D}$, $\mathcal{M}\left(\hat{s}_{1}(\sigma,s,t),\hat{s}_{2}^{(\lambda)}(\sigma,s,t)\right)$ still vanishes at large $\sigma$. In general, choosing arbitrary $\hat{s}_{1},\hat{s}_{2}$ functions can map the singularities along the contour $\mathcal{C}_{1}$ to very complicated contours. In order to have the simplest analytic structure in the $\sigma$ plane, we will restrict to functions that map straight line contours to straight line contours and satisfy the conditions we require. The only allowed class of mappings are affine linear transformations for $\hat{s}_{1}$ and fractional linear transformations for $\hat{s}_{2}$. Therefore, we consider
$\hat{s}_{1}(\sigma,s,t)=a\sigma+b,\quad\hat{s}_{2}^{\lambda}(\sigma,s,t)=\frac{-\lambda\sigma+b}{\sigma+d}$ (2.6)
Now our goal is to obtain an $s\leftrightarrow t$ symmetric dispersion relation. So we further demand
$\hat{s}_{1}(\sigma,s,t)|_{\sigma=s,t}=s,t,\quad\hat{s}_{2}^{(\lambda)}(\sigma,s,t)|_{\sigma=s,t}=t,s$ (2.7)
This uniquely fixes $\hat{s}_{1}(\sigma,s,t)=\sigma$ and $\hat{s}_{2}^{(\lambda)}(\sigma,s,t)=\frac{(s+\lambda)(t+\lambda)}{\sigma+\lambda}-\lambda$. We see that $\lambda$ is left over as a free parameter. Now let us explain the form of the kernel in (2.2). The first two terms are obvious from the requirement of crossing symmetry. The third term follows from the fact that $\hat{s}_{2}^{(\lambda)}(\sigma,s,t)|=t,s$ when $\lambda=-s,-t$. This tells us that the parametric dispersive representation must turn into the fixed-$t,s$ dispersion relations in the limits when $\lambda=-s,-t$, and therefore, we must necessarily have the third term in the kernel to cancel the poles in $s,t$-channels when $\lambda=-s,-t$.
where $\Gamma$ is a large circle at infinity. From here, the parametric dispersion relation in (2.1) follows similarly to the fixed-$t$ case by shrinking the contour and picking up the singularities from the kernel and the amplitude. Note that while it may seem like there are new singularities being introduced at $\sigma+\lambda=0$ but this is avoided because $\mathcal{M}\left(\sigma,\frac{(s+\lambda)(t+\lambda)}{\sigma+\lambda}-\lambda\right)\Big|_{\sigma=-\lambda}=\mathcal{M}\left(-\lambda,\infty\right)=\mathcal{M}\left(\infty,-\lambda\right)=0$ due to our assumed large $s$ behavior.
While the above method is nice and intuitive, we will now present a different way to derive the parametric dispersion relation which rigorously shows that the form of $\hat{s}_{2}^{(\lambda)}(\sigma,s,t)$ has to be as we saw above once we fix the kernel to be of the form in (2.2). This method also generalises easily to the case of three-channel crossing symmetric amplitudes, which we discuss later. Let us treat $\hat{s}_{2}^{(\lambda)}(\sigma,s,t)$ as an arbitrary, as-of-now undetermined function We only assume for now that $\lim_{|\sigma|\to\infty}\hat{s}_{2}^{(\lambda)}(\sigma,s,t)\in\mathcal{D}$. Then
@@
and we can write (2.8). If the amplitude grows faster at large $\sigma$ than we have assumed, we can introduce additional counter-terms/subtraction-terms in the integrand to ensure that the integral vanishes. We discuss such cases in the next section. Now, let us contract the contour. We pick up singularities from both the kernel and the amplitude. The contributions from the amplitude lead to the following.
@@
where $\mathcal{A}^{(s)}(s,t)$ and $\mathcal{A}^{(t)}(s,t)$ are the discontinuities of the amplitude in the $s$-channel and $t$-channel respectively along the contour $\mathcal{C}_{1}$. Due to $s\leftrightarrow t$ symmetry of the amplitude, $\mathcal{A}^{(t)}(s,t)=\mathcal{A}^{(s)}(t,s)$. Crossing symmetry dictates that $\hat{s}_{2}^{(\lambda)}(\sigma,s,t)$ must map $\mathcal{C}_{2}$ to $\mathcal{C}_{1}$. We now demand that the two terms in (2.10) be equal. This can be achieved by doing a coordinate transformation $\sigma=\sigma(\tau)$ in the second term as follows
$\frac{1}{\pi}\int\mathrm{d}\tau\frac{d\sigma}{d\tau}\mathcal{H}^{(\lambda)}(\sigma(\tau),s,t)\mathcal{A}^{(t)}(\sigma(\tau),\hat{s}_{2}^{(\lambda)}(\sigma(\tau),s,t))=\frac{1}{\pi}\int\mathrm{d}\tau\frac{d\sigma}{d\tau}\mathcal{H}^{(\lambda)}(\sigma(\tau),s,t)\mathcal{A}^{(s)}(\hat{s}_{2}^{(\lambda)}(\sigma(\tau),s,t),\sigma(\tau))\,$, (2.11)
and demanding that the following conditions be met
@@
$\frac{d\sigma}{d\tau}\frac{d}{d\sigma}\ln\frac{(\sigma-s)(\sigma-t)}{(\lambda+\sigma)}=\frac{d}{d\tau}\ln\frac{(\tau-s)(\tau-t)}{(\lambda+\tau)}\,$, (2.13)
@@
where $c$ is an integration constant. The second condition fixes $c=1$ and also maps $\mathcal{C}_{2}$ to $\mathcal{C}_{1}$. We thus get a quadratic equation for $\tau$ with two solutions:
@@
The first solution has no $s,t$ dependence and therefore we discard it. The second solution leads to precisely the form of $\hat{s}_{2}^{(\lambda)}(\sigma,s,t)$ that we found before. Using the second solution, the contribution from the singularities of the amplitude is
@@
Now, we come to the singularities that arise from the kernel $\mathcal{H}(s,t,\lambda)$. As explained before, the apparent singular term at $\sigma=-\lambda$ is actually not singular due to our assumed large $\sigma$ behaviour. Therefore,
@@
Figure 2: A typical depiction of the singularities in the integrand in the complex $\sigma$ plane. We consider the amplitude $\frac{1}{\sqrt{2-s}}+\frac{1}{\sqrt{2-t}}$. Then $\mathcal{C}_{1}$ is the contour $\sigma\geq 2$ and $\mathcal{C}_{2}$ is the contour $\hat{s}_{2}^{(\lambda)}(\sigma,s,t)\geq 2$. In the plot, we choose $s=1+i,\>t=2+2i,\>\lambda=1$ to determine $\mathcal{C}_{2}$. It is the line joining $(1/3,7/3)$ and (1,0)(-1,0).
@@ 3.1 Derivation
when $\lim_{|\sigma|\to\infty}\hat{s}_{2}^{(\lambda)}(\sigma,s_{1},s_{2},s_{3})\in\mathcal{D}$. $\hat{s}_{2}^{(\lambda)}(\sigma,s_{1},s_{2},s_{3})$ is an as-of-yet undetermined function. Now, the following contour integral along a large circle at infinity vanishes.
@@ 4.3 One-loop box diagram
Dispersion relations can be used in the evaluation of Feynman diagrams [33, 34]. The one-loop box diagram in four dimensions with two massive and two massless propagators takes the following form [35]
@@ 5.4 Bounds on weakly-coupled gravitational EFTs
Figure 8: Bounds with gravity for dilaton-like amplitudes in $d=5$. The region to the right of the curves is allowed. As can be seen, the bounds converge as $k_{max}$ is increased. We define $W_{1,0}=-\dfrac{2g_{2}}{8\pi G}$ and $W_{0,1}=\dfrac{g_{3}}{8\pi G}$ to match the conventions in [49] and aid comparison. Note that the bounds here are stronger compared to [49] because we only a consider a restricted subspace of meromorphic amplitudes.